# Advantages of fixing spins in quantum annealing

### v1

@@ Abstract
Many quantum-classical methods using QA have been proposed to address large combinatorial optimization problems [26, 29, 23, 25, 27]. To clarify the effect of fixing spins as a size-reduction method used in most of them, we parameterized the methods of fixing spins and investigated the properties.
@@
where $\mathcal{H}_{\mathrm{p}}^{\prime}$ is subproblem Hamiltonian and $\mathcal{H}_{\mathrm{d}}^{\prime}$ is driver Hamiltonian that is transverse magnetic fields for $n$ spins. Subproblem Hamiltonian is represented in the following form:
@@
Figure 1: (Color online) (a) Number of spins in the subproblem $n$ dependence of mean values of the residual energy $\Delta E_{\mathrm{res}}$ for the $100$ instances with $N=168$. Dashed lines indicate the residual energy $E^{\mathrm{S}}_{\mathrm{ref}}$ derived by only pre-processing. The lines linearly interpolate the plots of the experimental results. (b) Mean value of the residual energy $\Delta E_{\mathrm{res}}^{\mathrm{S}}$ for extracted spins at each $n$ for the $100$ instances with $N=168$. Dashed line shows the value of zero if the QA obtains reference energy. In both graphs, error rates are $p_{\mathrm{err}}=0.0$, $p_{\mathrm{err}}=0.1$, and $p_{\mathrm{err}}=0.5$, respectively. Error bars are the standard deviations in $100$ instances. QA time $\tau$ ranges $\tau=1~{}\mu\mathrm{s}$ from $\tau=2000~{}\mu\mathrm{s}$. When $n=168$, the problems are directly input into the quantum annealer. Solid lines between points are merely a guide to the eye.
@@
We consider the physical quantities described to investigate the properties of fixing spins. The energy of reference solutions, the energy obtained by HQA, and the subproblem’s reference solution energy are denoted by $E_{\mathrm{ref}}$, $E_{\mathrm{HQA}}$, and $E^{\mathrm{S}}_{\mathrm{ref}}$, respectively. $E^{\mathrm{S}}_{\mathrm{ref}}$ is also obtained by solving in Fixstars Amplify AE. We evaluated the performance of HQA by using the residual energy $\Delta E_{\mathrm{res}}\coloneqq E_{\mathrm{HQA}}-E_{\mathrm{ref}}$ and evaluated the performance of the QA part after fixing spins by using the subproblem’s residual energy $\Delta E_{\mathrm{res}}^{\mathrm{S}}\coloneqq E_{\mathrm{HQA}}-E^{\mathrm{S}}_{\mathrm{ref}}$. The minimum energy gap $\Delta_{\min}$ during the QA in the subproblem is used to evaluate the time enough to follow the ground state in the dynamic process of the QA.
@@
Figure 1 (a) shows the $n$ dependence of $\Delta E_{\mathrm{res}}$ for some $p_{\mathrm{err}}$ and $\tau$. The blue points indicate the case of adopting fixing spins with $p_{\mathrm{err}}=0$. In this case, increasing $n$ renders $\Delta E_{\mathrm{res}}$ large. This phenomenon could be attributed to the large system requiring longer annealing time to obtain the ground state of the subproblem. These results indicate an upper bound of the number of spins exists in the subproblem, which depends on $\tau$ to obtain the optimal solution in a quantum annealer. In cases of erroneous fixing spins (purple and brown points in Fig. 1 (a)), the results indicate that the optimal number $n^{\mathrm{opt}}$ of spins exists to obtain less $\Delta E_{\mathrm{res}}$. When $p_{\mathrm{err}}=0.1$, the nearest number with the optimal number $n^{\mathrm{opt}}$ is $n=96$ for $\tau=1~{}\mu\mathrm{s}$, $n=128$ for $\tau=10~{}\mu\mathrm{s}$, $n=148$ for $\tau=100~{}\mu\mathrm{s}$, $n=148$ for $\tau=1000~{}\mu\mathrm{s}$ and $n=148$ for $\tau=2000~{}\mu\mathrm{s}$. When $p_{\mathrm{err}}=0.1$, which is a small error rate, the system reaches the ground state of the subproblem, increasing $n$, rendering the residual energy small. After exceeding the optimal number $n^{\mathrm{opt}}$, increasing $n$ makes $\Delta E_{\mathrm{res}}$ large, that is, the system cannot reach the ground state of the subproblem. This phenomenon indicates a trade-off between the classical method’s solution accuracy and the QA’s solution accuracy. When $p_{\mathrm{err}}=0.5$, which is a large error rate, because the accuracy of the classical method is too worse than QA, the effects of increasing $n$ are dominant. Hence, increasing $n$ is an excellent way to obtain less residual energy when the accuracy of the classical method is worse than a certain standard.
@@
Figure 1 (b) reveals the $n$ dependence with the residual energy $\Delta E_{\mathrm{res}}^{\mathrm{S}}$ of the subproblem for some $\tau$ and $p_{\mathrm{err}}$. The annealing time comparison reveals that HQA performs well when the annealing time is long, and the annealing time at which $\Delta E_{\mathrm{res}}^{\mathrm{S}}$ begins to become nonzero differs for each $n$. These results indicate that $n^{\mathrm{opt}}$ depends on $\tau$. The error rate comparisons reveal that $n$ at $\Delta E_{\mathrm{res}}^{\mathrm{S}}=0$ varies approximately according to $p_{\mathrm{err}}$. Therefore, the error rate affects the performance of QA as well as HQA performance. These results show that the appropriate number of spins exists and depends on $\tau$ and $p_{\mathrm{err}}$ to obtain the best solution among solutions derived by the same classical method and QA. For all cases except for adopting fixing spins with $p_{\mathrm{err}}=0$, the lowest residual energy among energies derived by HQA or only QA is smaller than residual energy derived by only the classical method. This result indicates that using a quantum-classical hybrid method is an effective way to obtain a better solution unless the classical method obtains the ground state.
Reducing the number of spins leads to reducing the search space. The solution accuracy improves as the number of spins is smaller in the actual hardware, as shown [37]. The method of fixing spins is one of the causes of improvement in solution accuracy. Expanding the $\Delta_{\mathrm{min}}$ is also the cause of the improvement. Figure 2 shows $\Delta_{\mathrm{min}}$ scaling. These show that $\Delta_{\mathrm{min}}$ expands as the fixed spins increase. The expansion degree depends on the subproblem’s number of spins. $\mathcal{H}^{\prime}$ (in Eq. (5)) has larger $\Delta_{\min}$ than $\mathcal{H}$ (in Eq. (1)) and $\Delta_{\min}$ decreases as $n$ increases. From the viewpoint of the adiabatic theorem shown in Eq. (4), as $\Delta_{\min}$ becomes smaller, long $\tau$ is required to obtain the ground state. Since $\Delta_{\min}$ tends to be increased as increasing the number of fixed spins, fixing spins has effects on QA to shorten $\tau$ to obtain the ground state. Therefore, the probability of obtaining the ground state increases when $n$ is small. The performance of QA for the subproblem is considered to increase by fixing spins. In Fig. 2, using erroneous fixing spins, $\Delta_{\min}$ also becomes smaller as $n$ increases. However, when using erroneous fixing spins, the degree of the expansion of $\Delta_{\mathrm{min}}$ is smaller than using fixing spins with no errors. The other $\Delta_{\mathrm{min}}$ dependence of $p_{\mathrm{err}}$ is not confirmed because $N$ is small in this numerical analysis.
@@
Figure 2: (Color online) Mean value of the minimum energy gap $\Delta_{\mathrm{min}}$ at each $n$ for different $100$ instances with $N=12$. Error bars are the standard deviations in $100$ instances. Dashed line shows the minimum energy gap without fixing spins. Solid lines between points are merely a guide to the eye.

### v2

@@ Abstract
Many quantum-classical methods using QA have been proposed to address large combinatorial optimization problems [26, 24, 23, 25, 27]. These quantum-classical approaches surpass the performance of both purely classical and purely quantum methods. To clarify the effect of fixing spins as a size-reduction method used in most of them, we parameterized the methods of fixing spins and investigated the properties.
@@
where $\mathcal{H}_{\mathrm{p}}^{\prime}$ is subproblem Hamiltonian and $\mathcal{H}_{\mathrm{d}}^{\prime}$ is driver Hamiltonian that is transverse magnetic fields for $n$ spins. The spin indices $i$ in Eq. (2) are relabeled into $i^{\prime}$ based on the spins in subproblem Hamiltonian. Subproblem Hamiltonian is represented in the following form:
@@
Figure 1: (Color online) (a)-(c) Number of spins in the subproblem $n$ dependence of mean values of the residual energy $\Delta E_{\mathrm{res}}$ for the $100$ instances with $N=168$. Dashed lines indicate the residual energy $E_{\mathrm{ref}}(n)$ derived by only pre-processing. The lines linearly interpolate the plots of the experimental results. (d) Mean value of the residual energy $\Delta E_{\mathrm{res}}(n)$ for extracted spins at each $n$ for the $100$ instances with $N=168$. Dashed line shows the value of zero if the QA obtains reference energy. Each panel indicates (a) $p_{\mathrm{err}}=0.0$, (b) $p_{\mathrm{err}}=0.1$ and (c) $p_{\mathrm{err}}=0.5$, respectively. QA time $\tau$ ranges $\tau=1~{}\mu\mathrm{s}$ from $\tau=2000~{}\mu\mathrm{s}$. When $n=168$, the problems are directly input into the quantum annealer. Error bars in all figures are omitted for clarity since their sizes are smaller than the symbol size.
@@
We consider the physical quantities described to investigate the properties of fixing spins. The energy of reference solutions, the energy obtained by HQA, and the subproblem’s reference solution energy are denoted by $E_{\mathrm{ref}}$, $E_{\mathrm{HQA}}$, and $E_{\mathrm{ref}}(n)$, respectively. $E_{\mathrm{ref}}(n)$ is also obtained by solving in Fixstars Amplify AE. We evaluated the performance of HQA by using the residual energy $\Delta E_{\mathrm{res}}\coloneqq E_{\mathrm{HQA}}-E_{\mathrm{ref}}$ and evaluated the performance of the QA part after fixing spins by using the subproblem’s residual energy $\Delta E_{\mathrm{res}}(n)\coloneqq E_{\mathrm{HQA}}-E_{\mathrm{ref}}(n)$. The minimum energy gap $\Delta_{\min}$ during the QA in the subproblem is used to evaluate the time enough to follow the ground state in the dynamic process of the QA.
@@
Figure 1 (a)-(c) shows the $n$ dependence of $\Delta E_{\mathrm{res}}$ for some $p_{\mathrm{err}}$ and $\tau$. Figure 1 (a) indicates the case of adopting fixing spins with $p_{\mathrm{err}}=0$. In this case, increasing $n$ renders $\Delta E_{\mathrm{res}}$ large. This phenomenon could be attributed to the large system requiring longer annealing time to obtain the ground state of the subproblem. These results indicate an upper bound of the number of spins exists in the subproblem, which depends on $\tau$ to obtain the optimal solution in a quantum annealer. In cases of erroneous fixing spins (in Fig. 1 (b) and (c)), the results indicate that the optimal number $n^{\mathrm{opt}}$ of spins exists to obtain less $\Delta E_{\mathrm{res}}$. When $p_{\mathrm{err}}=0.1$, the nearest number with the optimal number $n^{\mathrm{opt}}$ is $n=96$ for $\tau=1~{}\mu\mathrm{s}$, $n=128$ for $\tau=10~{}\mu\mathrm{s}$, $n=148$ for $\tau=100~{}\mu\mathrm{s}$, $n=148$ for $\tau=1000~{}\mu\mathrm{s}$ and $n=148$ for $\tau=2000~{}\mu\mathrm{s}$. When $p_{\mathrm{err}}=0.1$, which is a small error rate, the system reaches the ground state of the subproblem, increasing $n$, rendering the residual energy small. After exceeding the optimal number $n^{\mathrm{opt}}$, increasing $n$ makes $\Delta E_{\mathrm{res}}$ large, that is, the system cannot reach the ground state of the subproblem. This phenomenon indicates a trade-off between the classical method’s solution accuracy and the QA’s solution accuracy. When $p_{\mathrm{err}}=0.5$, which is a large error rate, because the accuracy of the classical method is too worse than QA, the effects of increasing $n$ are dominant. Hence, increasing $n$ is an excellent way to obtain less residual energy when the accuracy of the classical method is worse than a certain standard.
@@
Figure 1 (d) reveals the $n$ dependence with the residual energy $\Delta E_{\mathrm{res}}(n)$ of the subproblem for some $\tau$ and $p_{\mathrm{err}}$. The annealing time comparison reveals that HQA performs well when the annealing time is long, and the annealing time at which $\Delta E_{\mathrm{res}}(n)$ begins to become nonzero differs for each $n$. These results indicate that $n^{\mathrm{opt}}$ depends on $\tau$. The error rate comparisons reveal that $n$ at $\Delta E_{\mathrm{res}}(n)=0$ varies approximately according to $p_{\mathrm{err}}$. Therefore, the error rate affects the performance of QA as well as HQA performance. These results show that the appropriate number of spins exists and depends on $\tau$ and $p_{\mathrm{err}}$ to obtain the best solution among solutions derived by the same classical method and QA. For all cases except for adopting fixing spins with $p_{\mathrm{err}}=0$, the lowest residual energy among energies derived by HQA or only QA is smaller than residual energy derived by only the classical method. This result indicates that using a quantum-classical hybrid method is an effective way to obtain a better solution unless the classical method obtains the ground state.
@@
Figure 2: (Color online) Mean value of the minimum energy gap $\Delta_{\mathrm{min}}$ at each $n$ for different $100$ instances with $N=12$. Dashed line shows the minimum energy gap without fixing spins. Error bars are omitted for clarity since their sizes are smaller than the symbol size.
We conducted numerical simulation to investigate the dynamical properties of QA after fixing spins. Figure 2 shows $\Delta_{\mathrm{min}}$ scaling. These show that $\Delta_{\mathrm{min}}$ expands as the fixed spins increase. The expansion degree depends on the subproblem’s number of spins. $\mathcal{H}^{\prime}$ (in Eq. (5)) has larger $\Delta_{\min}$ than $\mathcal{H}$ (in Eq. (1)) and $\Delta_{\min}$ decreases as $n$ increases. From the viewpoint of the adiabatic theorem shown in Eq. (4), as $\Delta_{\min}$ becomes smaller, long $\tau$ is required to obtain the ground state. Since $\Delta_{\min}$ tends to be increased as increasing the number of fixed spins, fixing spins has effects on QA to shorten $\tau$ to obtain the ground state. Therefore, the probability of obtaining the ground state increases when $n$ is small. The performance of QA for the subproblem is considered to increase by fixing spins. In Fig. 2, using erroneous fixing spins, $\Delta_{\min}$ also becomes smaller as $n$ increases. However, when using erroneous fixing spins, the degree of the expansion of $\Delta_{\mathrm{min}}$ is smaller than using fixing spins with no errors. The other $\Delta_{\mathrm{min}}$ dependence of $p_{\mathrm{err}}$ is not confirmed because $N$ is small in this numerical analysis. Reducing the number of spins leads to reducing the search space. The solution accuracy improves as the number of spins is smaller in the actual hardware, as shown [36]. The fixing spins method which reduces the search space is one of the causes of improvement in solution accuracy. Also, these results on numerical simulation suggests that expanding the $\Delta_{\mathrm{min}}$ is also the cause of the improvement. Therefore, the energy gap properties explain the advantage of fixing spins obtained in previous studies[23, 24, 25, 26, 27, 28].